6. If 3x – 3 5 f(x) < x2 – 3x + 6 for x 2 0, find lim f(x). X-3
Answer #1

34-3 = f(x) 드 2²-3x + 6 take limit R tends to 3 lim (Bx-3) a lim f(x) = lim (22-3x+6) N> 3 N73 3(3)-3 { lim f(x) & 8 (3²-3(3)